Angles A and B are complements, and the measure of angle A is 8 times the measure of angle B. Find the measure(in degrees) of the supplement of A.
PLEASEE HELP && THANK YOU (:

Answer by checkley77(12569)

A=8B OR A-8B=0
A+B=90
A-8B=0 SUBTRACT.
------------------
9B=90
B=90/9
B=10 ANS.
A=8*10=80 ANS.